# The non-existence of some Galois representations of moderate degree in small characteristic

### v1

@@ Abstract.
We show unconditionally that the image of an irreducible mod $2$ representation $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$ that is unramified outside $2$ must be large. Under GRH, we then deduce the non-existence of such representations.
@@ 1. Introduction
Let $K$ be a number field, $d$ a positive integer, and $p$ a prime. The non-existence of $d$-dimensional mod $p$ Galois representations of $G_{K}:={\rm Gal}(\overline{K}/K)$ with small ramification plays an important role in proving Serre’s modularity conjecture (cf. [KW] for odd Galois representations with $K=\mathbb{Q}$ and $d=2$).
We will call a Galois representation $p$-ramified if it is unramified outside $\{p\}$. The first results on $p$-ramified Galois representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{2}(\overline{{\mathbb{F}}}_{p})$ are due to Tate [Ta] for $p=2$ and Serre [Se1] for $p=3$, followed by several authors who attacked this kind of problem for small $p,d$, and $K$ with $K:\mathbb{Q}]\leqslant 2$ ([Mo2],[MT],[Sen], [GM]).
@@
Assume GRH.
There exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{2})$ with $d\in\{2,3,4\}$.
There exist no irreducible totally real $3$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{3})$ with $d\in\{2,3,4\}$.
There exist no irreducible totally real $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{2})$ with $d\in\{5,6,7,8\}$.
Note that the case $p=2$, $d\leqslant 4$ was proved in [Mo2] by Moon without using GRH, but under the assumption that $\mathbb{Q}_{\overline{\rho}}:=\overline{{\mathbb{Q}}}^{{\rm Ker}(\overline{\rho})}$ is totally real.
@@
Assume GRH. There exist no self-dual $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{5}(\overline{{\mathbb{F}}}_{2})$ of length $2$.
Note that “of length 22” is used here in the representation-theoretic sense, namely that the Galois module has an irreducible submodule and the quotient by it is irreducible. (This is not to be confused with Moon’s notion of $p$-length, which is about finite groups and plays an essential role in the proof of Theorem A.) We also remark that for each subfield $F$ of $\overline{{\mathbb{F}}}_{2}$, any $2$-ramified self-dual Galois representation $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{5}(F)$ contains the trivial representation as a direct factor, so “of length 22” is the best we can hope for.
The second flavor of result we present is unconditional and focuses on symplectic representations, starting with the ${\rm GSp}_{4}(\mathbb{F}_{2})$-valued case:
@@
There exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\mathbb{F}_{2})$.
The proof uses the isomorphism ${\rm GSp}_{4}(\mathbb{F}_{2})\simeq S_{6}$ and associates to $\overline{\rho}$ a rational polynomial of degree at most $6$, approach that may also be of use in the study of Galois representations coming from $2$-torsion points on abelian surfaces, as in [GY].
In the more general ${\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$-valued case we show:
@@
Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$ be an irreducible $2$-ramified representation. Then ${\rm Im}(\overline{\rho})$ is either isomorphic to the Suzuki group ${\rm Sz}(\mathbb{F}_{2^{r}})$ for some odd $r>1$, or is equal to ${\rm Sp}_{4}(\mathbb{F}_{2^{s}})$ for some $s>1$.
In particular, there exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$ such that $|{\rm Im}(\overline{\rho})|<29120$.
We note that the proof may be applicable to representations of $G_{K}$ for number fields $K\neq{\bf Q}$. The key ingredients are: the classification of maximal subgroups inside ${\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$, the non-existence results for $2$-dimensional representations over certain quadratic fields, and the Jones–Roberts database of number fields [JR]. Eventually, the problem is reduced to the case of Theorem C.
@@
Assume GRH. There exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GSp}_{4}(\overline{{\mathbb{F}}}_{2})$.
A crucial ingredient, used in many places in this paper as well as in several of the key references, is provided by the discriminant bounds of Odlyzko, Poitou, and Serre, and the related tables of discriminant bounds computed by Odlyzko, see [Od]. This is also where the dependency on GRH stems from.
This paper is organized as follows. After some preparatory work in Section 2, we build on Moon’s methods from [Mo1, Mo2] in Section 3 to prove Theorem A, then obtain Corollary B as a consequence in Section 4. In Section 5 we consider ${\rm GSp}4(\mathbb{F}_{2})$-valued representations and prove Theorem C, and in Section 6 we extend this to obtain a proof of Theorem D and Corollary E. The appendix is devoted to detailing the Magma ([BCP]) and SageMath ([Sage]) computations used in this paper; we hope this will facilitate reproducibility.
@@ 2.2. Modular representations of some finite groups
The same argument and conclusion apply to the case of irreducible $\overline{{\mathbb{F}}}_{p}$-representations of the dihedral group $D_{n}=(\mathbb{Z}/n\mathbb{Z})\rtimes(\mathbb{Z}/2\mathbb{Z})$ for any odd integer $n$ divisible by $p$.
@@
    There are no irreducible Galois representations $\overline{\rho}\colon G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{p})$ such that the Galois group of $\mathbb{Q}_{\overline{\rho}}$ is a $p$-group.
@@
    For $p$ odd, there are no $p$-ramified irreducible Galois representations $\overline{\rho}\colon G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{p})$ such that the Galois group of $\mathbb{Q}_{\overline{\rho}}$ is isomorphic to $(\mathbb{Z}/p\mathbb{Z})^{m}\rtimes(\mathbb{Z}/2\mathbb{Z})$.
@@ 2.3. Discriminant upper bounds from $p$-Sylow structure of Galois groups
Let $F/\mathbb{Q}_{p}$ be a finite unramified extension. Let $E/F$ be a totally ramified extension with Galois group $I$. Suppose $I$ has $p$-length NN and fix a series of the form (1) for the $p$-Sylow subgroup $P$ of $I$. Let $e_{0}=[I\colon P]$, let $r$ be the remainder of the division of $e_{0}$ by p1p-1, and let $e_{N}$ be the order of $I$. Then the different of $E/F$ divides $p^{C}$, where
@@
Here is the strategy we are going to employ: a priori we have no control over the quantity in the parenthesis in Proposition 2.7, so we simply use C<N+1+N/(p1)C<N+1+N/(p-1). This gives us a bound on $|d_{K}|^{1/n}$, hopefully small enough to appear in Odlyzko’s tables. The latter give us an upper bound on $n$, say $n\leqslant n_{\text{max}}$. At this point we return to the expression in Proposition 2.7 and minimize the quantity
@@
3. Non-existence of ${\rm GL}_{d}(\overline{{\mathbb{F}}}_{p})$-valued representations (Thm A)
@@
Given an irreducible $p$-ramified Galois representation $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{p})$, we let $K$ be the field fixed by the kernel of $\overline{\rho}$ and $G$ be the Galois group of $K/\mathbb{Q}$. We also let $n=[K:\mathbb{Q}]=|G|$.
@@
If $p\in\{2,3\}$, there exist no irreducible $p$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{p})$ with $\mathbb{Q}_{\overline{\rho}}=K$.
@@
With $N=0$ and $p=2$, we get from Proposition 2.7 that $|d_{K}|^{1/n}<2$, and then from Odlyzko’s (unconditional) Table 2 that $n<3$, which forces the Galois group to be abelian and hence not have irreducible representations of dimension >1>1.
@@
If $p\in\{2,3\}$, there exist no irreducible $p$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{p})$ with $\mathbb{Q}_{\overline{\rho}}=K$.
@@
This leaves as only possibility $n=72$, which gives us the tighter bound $C<\frac{5}{2}-\frac{13}{72}<2.320$. We deduce that $|d_{K}|^{1/n}<12.792$ and thus $n<40$, and we are done. ∎
@@
    if $p=2$, there exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{2})$ with $\mathbb{Q}_{\overline{\rho}}=K$.
@@
    if $p=3$, there exist no irreducible totally real $3$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{3})$ with $\mathbb{Q}_{\overline{\rho}}=K$.
@@
    At this point the set of candidates for $n$ is not quite empty! We are left with the case $n=18$, which we now rule out. There are three (isomorphism classes of) non-abelian groups of order $18$:
@@
Assuming GRH, there exist no irreducible totally real $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{2})$ with $\mathbb{Q}_{\overline{\rho}}=K$.
@@
Assume GRH.
There exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{2})$ with $d\in\{2,3,4\}$.
There exist no irreducible totally real $3$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{3})$ with $d\in\{2,3,4\}$.
There exist no irreducible totally real $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow{\rm GL}_{d}(\overline{{\mathbb{F}}}_{2})$ with $d\in\{5,6,7,8\}$.
@@
Follows from the previous Propositions, since the $p$-length of any finite subgroup of ${\rm GL}_{d}(\overline{{\mathbb{F}}}_{p})$ is $\leqslant\lceil\log_{2}(d)\rceil$, as shown at the beginning of [Mo1, Section 3]. ∎

### v2

# The non-existence of some Galois representations of moderate dimension in small characteristic

@@ Abstract.
We show unconditionally that the image of an irreducible mod $2$ representation $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$ that is unramified outside $2$ must be large. Under GRH, we then deduce the non-existence of such representations.
@@ 1. Introduction
Let $K$ be a number field, $d$ a positive integer, and $p$ a prime. The non-existence of $d$-dimensional mod $p$ Galois representations of $G_{K}:=\mathrm{Gal}(\overline{K}/K)$ with small ramification plays an important role in proving Serre’s modularity conjecture (cf. [KW] for odd Galois representations with $K=\mathbb{Q}$ and $d=2$).
We will call a Galois representation $p$-ramified if it is unramified outside $\{p\}$. The first results on $p$-ramified Galois representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\overline{\mathbb{F}}_{p})$ are due to Tate [Ta] for $p=2$ and Serre [Se1] for $p=3$, followed by several authors who attacked this kind of problem for small $p,d$, and $K$ with $K:\mathbb{Q}]\leqslant 2$ ([Mo2],[MT],[Sen], [GM]).
@@
Assume GRH. There exist no irreducible
  1. (a)
    22-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{2})$ with $d\in\{2,3,4\}$;
  2. (b)
    totally real $3$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{3})$ with $d\in\{2,3,4\}$;
  3. (c)
    totally real $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{2})$ with $d\in\{5,6,7,8\}$.
Note that the case $p=2$, $d\leqslant 4$ was proved in [Mo2] by Moon without using GRH, but under the assumption that $\mathbb{Q}_{\overline{\rho}}:=\overline{{\mathbb{Q}}}^{\ker(\overline{\rho})}$ is totally real.
@@
Assume GRH. There exist no self-dual $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{5}(\overline{\mathbb{F}}_{2})$ of length $2$.
Note that “of length 22” is used here in the representation-theoretic sense, namely that the Galois module has an irreducible submodule and the quotient by it is irreducible. (This is not to be confused with Moon’s notion of $p$-length, which is about finite groups and plays an essential role in the proof of Theorem A.) We also remark that for each subfield $F$ of $\overline{\mathbb{F}}_{2}$, any $2$-ramified self-dual Galois representation $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{5}(F)$ contains the trivial representation as a direct factor, so “of length 22” is the best we can hope for.
The second flavor of result we present is unconditional and focuses on symplectic representations, starting with the $\mathrm{GSp}_{4}(\mathbb{F}_{2})$-valued case:
@@
There exist no representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\mathbb{F}_{2})$ that are irreducible and $2$-ramified.
The proof uses the isomorphism $\mathrm{GSp}_{4}(\mathbb{F}_{2})\simeq S_{6}$ and associates to $\overline{\rho}$ a rational polynomial of degree at most $6$, approach that may also be of use in the study of Galois representations coming from $2$-torsion points on abelian surfaces, as in [GY].
In the more general $\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$-valued case we show:
@@
Let $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$ be an irreducible $2$-ramified representation. Then $\mathrm{Im}(\overline{\rho})$ is either isomorphic to the Suzuki group $\mathrm{Sz}(\mathbb{F}_{2^{r}})$ for some odd $r>1$, or is equal to $\mathrm{Sp}_{4}(\mathbb{F}_{2^{s}})$ for some $s>1$.
Hence there exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$ such that $|\mathrm{Im}(\overline{\rho})|<29120$.
We note that the proof may be applicable to representations of $G_{K}$ for number fields $K\neq\mathbb{Q}$. The key ingredients are: the classification of maximal subgroups inside $\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$, the non-existence results for $2$-dimensional representations over certain quadratic fields, and the Jones–Roberts database of number fields [JR]. Eventually, the problem is reduced to the case of Theorem C.
@@
Assume GRH. There exist no irreducible $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\overline{\mathbb{F}}_{2})$.
As an immediate consequence we have:
Corollary F.
Let $F$ be a holomorphic Siegel cusp form of degree $2$ and weight $(k_{1},k_{2})$ with $k_{1}\geqslant k_{2}\geqslant 1$ corresponding to $\big(\mathrm{Sym}^{k_{1}-k_{2}}\otimes\det^{k_{2}}\big)(\mathbb{C}^{2})$. Suppose that $F$ is a Hecke eigenform and let $\mathbb{Q}_{F}$ be the Hecke field of $F$ (the number field generated by the Hecke eigenvalues). Suppose also that the cuspidal automorphic representation $\pi_{F}$ attached to $F$ is unramified outside $2$. Let $v$ be a finite place of $\mathbb{Q}_{F}$ above $2$, let $\mathbb{F}_{v}$ be the residue field and $\overline{\rho}_{F,v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GSp}_{4}(\mathbb{F}_{v})$ be the semisimplification of the mod $v$ Galois representation attached to $F$. We have
  1. (1)
    if $\mathbb{F}_{v}=\mathbb{F}_{2}$, then $\overline{\rho}_{F,v}$ is trivial;
  2. (2)
    if we assume GRH, then $\overline{\rho}_{F,v}$ is trivial.
The existence of $\overline{\rho}_{F,v}$ follows from [Wei] when $k_{2}\geqslant 3$ and $F$ is neither a CAP form nor an endoscopic form. All Siegel holomorphic CAP and endoscopic forms for any weight have been classified (cf. [KWY, Sections 3.4 and 3.5] for a summary), and one can easily attach $\overline{\rho}_{F,v}$ from the shape of the spinor $L$-function of $F$. For the lower-weight case ($k_{2}\leqslant 2$), we may use the mod $2$ Hasse invariant to raise the weight and thereby apply the former results.
This paper is organized as follows. After some preparatory work in Section 2, we build on Moon’s methods from [Mo1, Mo2] in Section 3 to prove Theorem A, then obtain Corollary B as a consequence in Section 4. In Section 5 we consider $\mathrm{GSp}_{4}(\mathbb{F}_{2})$-valued representations and prove Theorem C, and in Section 6 we extend this to obtain a proof of Theorem D and Corollary E. The appendix is devoted to detailing the Magma ([BCP]) and SageMath ([Sage]) computations used in this paper; we hope this will facilitate reproducibility.
A crucial ingredient for proving the non-existence of representations, used in many places in this paper as well as in several of the key references, is provided by the discriminant bounds of Odlyzko, Poitou, and Serre, and the related tables of discriminant bounds computed by Odlyzko, see [Od]. This is also where the dependency on GRH stems from.
@@ 2.2. Modular representations of some finite groups
The same argument and conclusion apply to the case of irreducible $\overline{\mathbb{F}}_{p}$-representations of the dihedral group $D_{n}=(\mathbb{Z}/n\mathbb{Z})\rtimes(\mathbb{Z}/2\mathbb{Z})$ if $n$ is a power of the odd prime $p$.
@@
    There are no irreducible representations $\overline{\rho}\colon G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{p})$ such that the Galois group of $\mathbb{Q}_{\overline{\rho}}$ is a $p$-group.
@@
    For $p$ odd, there are no irreducible representations $\overline{\rho}\colon G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{p})$ that are $p$-ramified and such that $\mathrm{Gal}(\mathbb{Q}_{\overline{\rho}}/\mathbb{Q})\simeq(\mathbb{Z}/p\mathbb{Z})^{m}\rtimes(\mathbb{Z}/2\mathbb{Z})$.
@@ 2.3. Discriminant upper bounds from $p$-Sylow structure of Galois groups
Let $F/\mathbb{Q}_{p}$ be a finite unramified extension. Let $E/F$ be a totally ramified extension with Galois group $I$. Suppose $I$ has $p$-length NN and fix a series of the form (1) for the $p$-Sylow subgroup $P$ of $I$, with corresponding positive integers $m_{1},\dots,m_{N}$. Let $e_{0}=[I\colon P]$, let $r$ be the remainder of the division of $e_{0}$ by p1p-1, and let $e_{N}$ be the order of $I$. Then the different of $E/F$ divides $p^{C}$, where
@@
Here is the strategy we are going to employ: a priori we have no control over the quantity in the parenthesis in Proposition 2.7, so we simply use
C<N+1+N/(p1).C<N+1+N/(p-1).
This gives us a bound on $|d_{K}|^{1/n}$, hopefully small enough to appear in Odlyzko’s tables. The latter give us an upper bound on $n$, say $n\leqslant n_{\text{max}}$. At this point we return to the expression in Proposition 2.7 and minimize the quantity
@@
3. Non-existence of $\mathrm{GL}_{d}(\overline{\mathbb{F}}_{p})$-valued representations (Thm A)
@@
Given an irreducible $p$-ramified Galois representation $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{p})$, we let $K$ be the field fixed by the kernel of $\overline{\rho}$ and $G$ be the Galois group of $K/\mathbb{Q}$. We also let $n=[K:\mathbb{Q}]=|G|$.
@@
If $p\in\{2,3\}$, there exist no irreducible representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{p})$ with $\mathbb{Q}_{\overline{\rho}}=K$ and $p$-ramified.
@@
With $N=0$ and $p=2$, we get from Proposition 2.7 that $|d_{K}|^{1/n}<2$, and then from Odlyzko’s (unconditional) Table 2 that $n<3$, so the Galois group is abelian and hence has no irreducible representations of dimension >1>1.
@@
If $p\in\{2,3\}$, there exist no irreducible representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{p})$ with $\mathbb{Q}_{\overline{\rho}}=K$ and $p$-ramified.
@@
This leaves as only possibility $n=72$, which gives us the tighter bound
$C<\frac{5}{2}-\frac{13}{72}<2.320.$
We deduce that $|d_{K}|^{1/n}<12.792$ and thus $n<40$, and we are done. ∎
@@
    if $p=2$, there exist no irreducible representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{2})$ with $\mathbb{Q}_{\overline{\rho}}=K$ and $2$-ramified;
@@
    if $p=3$, there exist no irreducible representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{3})$ with $\mathbb{Q}_{\overline{\rho}}=K$ totally real and $3$-ramified.
@@
    At this point the set of candidates for $n$ is not yet empty! We are left with the case $n=18$, which we now rule out. There are three (isomorphism classes of) non-abelian groups of order $18$:
@@
Assuming GRH, there exist no irreducible totally real $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{2})$ with $\mathbb{Q}_{\overline{\rho}}=K$.
@@
Assume GRH. There exist no irreducible
  1. (a)
    22-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{2})$ with $d\in\{2,3,4\}$;
  2. (b)
    totally real $3$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{3})$ with $d\in\{2,3,4\}$;
  3. (c)
    totally real $2$-ramified representations $\overline{\rho}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{d}(\overline{\mathbb{F}}_{2})$ with $d\in\{5,6,7,8\}$.
@@
Follows from the previous Propositions, since the $p$-length of any finite subgroup of $\mathrm{GL}_{d}(\overline{\mathbb{F}}_{p})$ is $\leqslant\lceil\log_{2}(d)\rceil$, as shown at the beginning of [Mo1, Section 3]. ∎